1. Thanks to Sugiyama-kun's preparation, we were able to adjust the resonant frequency of H1 folded type accelerometer to the target value of 0.2Hz today.
2. Although the resonant frequency of H2 and H3 has not been adjusted yet, the LVDTs driver for the three accelerometers in ITMX suspension, including H1, have been adjusted. So we can tune the frequency when we have time.
1. H1 folded accelerometer
The resonant frequency of H1 could be adjusted by putting some weights on it as shown in Fig. 3. Please refer to Takahashi-san's handwritten log attched in Fig.2 and the Diaggui file in /user/Miyo/dropbox/LOG/210426/ for details. After the final adjustment, we measured the transfer function, but we did not optimize the dynamic range of the LVDT, so we spent the rest of the day adjusting the LVDT driver, as described below.
2. Adjustment of LVDT dynamic range
The rest of the day was spent optimizing the gain and phase of the LVDT driver (S1403051) so that the LVDT signal would not saturate the ADC at the maximum tilt of the accelerometer pendulum. The procedure is as follows. First, the accelerometer was released and tilted to one side or the other. At that time, the accelerometer tilted to a certain inclination and then settled down, so we adjusted the gain and phase of the LVDT so that the value at that time would not be saturated by the ADC. The adjustment was done as follows. First, we turned the gain adjustment knob to set the count to about 30000. After that, turn the phase adjustment knob so that the count value is maximized. If, after optimizing the phase, the count increases or decreases, turn the gain adjustment knob to 30000 counts again. When we could confirm that the LVDT signal is not saturated on one side, then tilt the accelerometer to the other side and confirm that it is not saturated. If it is saturated, adjust the gain to 30000 counts. In this way, we adjusted the LVDT signal not to be saturated when the accelerometer pendulum was tilted on both sides.
After the adjustment, the transfer function was measured again only for H1. As shown in Fig. 1, the resonance frequency could be adjusted to 0.20 Hz.
we will do the resonant frequency tuning work for H2 and H3